OpenStudy (anonymous):

How do I find the diameter of (x + 4)^2 + (y – 9)^2 = 18?

OpenStudy (phi):

the equation for a circle with a center at (h,k) and with radius r is \[ (x-h)^2 + (y-k)^2 = r^2 \] match that to your equation. what is r^2 in your equation? what is r in your equation then remember the diameter is 2 times the radius. i.e. d= 2r

OpenStudy (anonymous):

Which is h and which is k? @phi

OpenStudy (phi):

the h goes with the x and the k goes with the y. But in this case, all you care about is "what is r^2" ?

OpenStudy (anonymous):

r^2 refers to the radius. Isn't it 6√2 or 8.485?

OpenStudy (anonymous):

Wait no.. √18 or 3√2

OpenStudy (phi):

r^2 is 18 (one step at a time) r is the square root of 18 \[ r= \sqrt{18} = \sqrt{2\cdot 9} = \sqrt{9} \cdot \sqrt{2} = 3\sqrt{2} \] last step, the diameter is 2 times the radius

OpenStudy (anonymous):

Right, got it. So 3√2*2=6√2.

OpenStudy (phi):

yes, 6√2 is the diameter of your circle.

OpenStudy (anonymous):

Okay, thank you so much! I really appreciate it. c:

OpenStudy (phi):

yw